OpenStudy (anonymous):

Find the derivative of f(x) = -7/x at x = -3.

OpenStudy (anonymous):

okay wait

OpenStudy (anonymous):

\[-7(\frac{ 1 }{ x })\] \[f(x)=-7x ^{-1}\] \[7x ^{1-1}\] =\[f \prime(x)=7\] i get that !! :\

hartnn (hartnn):

actually, \(\Large \dfrac{d}{dx}x^n = n x^{n-1}\) so it won't be 7

OpenStudy (anonymous):

\[7x ^{-1-1} = \frac{ 7 }{ x ^{2} }\]

hartnn (hartnn):

^ and then plug in x = -3 there

OpenStudy (anonymous):

ahaaaa okay so it will be \[f \prime(x)=7x ^{-2}\]

OpenStudy (anonymous):

my answer choices are 3/7 9/7 7/9 7/3

OpenStudy (anonymous):

then put x=-3 it will be \[7(-3)^{-2}=\frac{ 7 }{ 9 } =0.77\]

OpenStudy (anonymous):

thank you sooo much!!